OpenStudy (anonymous):

Find the constant term that must be added to the expression to make it a perfect square trinomial.

OpenStudy (anonymous):

x^2-5/2x

OpenStudy (caozeyuan):

25/16

OpenStudy (cwrw238):

(-5/2 ) / 2 = -5/ 4 (-5/4)^2 = 25/16

OpenStudy (jhannybean):

This is part of competing the square. You want to pput your equation in the form of \(ax^2 +bx+c\) and you're given \(ax^2+bx\). In order to find your "c" value, you take b,divide it by 2, and square it. The formula in order to find c follows the format \[\large c= \left(\frac{b}{2}\right)^2\]\[\large(x^2 -\frac52x)\]\[\large \left(-\frac{\frac{5}{2}}{2}\right)^2=\left(-\frac{5}{4}\right)^2=\frac{25}{16}\]\[\large \left( x^2-\frac52 x+\frac{25}{16} \right)-\frac{25}{16}\]Completed square. \[\large \left( x-\frac{5}{4}\right)^2 -\frac{25}{16}\]